Assume a Poisson distribution with λ=4.8.

Find the following probabilities

a. X=1 b.

X<1     

c.

X>1.   

d.

X≤1.   
Answer #1

a) P(X = 1)

= 0.0395

b) P(X < 1)

= P(X = 0)

= 0.0082

c) P(X > 1)

= 1 - P(X < 1)

= 1 - [P(X = 0) + P(X = 1)]

= 1 - [0.0395 + 0.0082]

= 0.9523

d) P(X < 1)

= P(X = 0) + P(X = 1)

= 0.0395 + 0.0082

= 0.0477